# CAAP: Context-Aware Action Planning Prompting to Solve Computer Tasks with Front-End UI Only

### v2

@@ 5.2 WebShop
WebShop (Yao et al., 2022b) is a simulated e-commerce web environment designed to evaluate agents’ performance on tasks that more closely resemble real-world scenarios. To successfully complete WebShop tasks, the agent must understand textual instructions and requirements for the target product and navigate various types of web pages to locate it, out of a pool of 1.18 million real products scraped from amazon.com. The environment provides a total of 12,087 instructions, divided into 10,587 for training, 1,000 for development, and 500 for testing. Among them, we utilized only the 500 test instructions for evaluation purposes.
Models and datasets
The web browser for the WebShop environment can be resized freely, and for convenience, we used a fixed resolution of 1276 × 1153 pixels for both data collection and evaluation. Unlike MiniWoB++, the dataset used to train the visual observer for WebShop was generated entirely through an automated process. Given an instruction, a simple helper agent we created extracted a search term, performed a search, and clicked the first item in the results. Screen images were captured throughout this process, and annotation data was collected using Selenium for each image, resulting in a total of 300 annotated samples. The decision-making model, supported action types, and function-calling approach are entirely consistent with those employed in the MiniWoB++ experiments. For the evaluation of the CAAP agent, a single human demonstration is applied universally to all instructions.
Experiment result
@@
There has been limited research on constructing agents that rely solely on image inputs, and Pix2Act is the only method whose performance has been validated on the WebShop benchmark. Table 2 presents a comparison between our approach and Pix2Act on WebShop. (Comparisons with other agent-based methods that rely on additional HTML input, on top of image input, are presented in Section 6 of Appendix B.2.) While Pix2Act employed 1,012 human demonstrations to train its agent model, our approach relied on just a single human demonstration, along with 300 samples collected through a fully automated process for training the visual observer module. Despite the significantly reduced manual data collection effort, our agent achieved a task score of 62.3, a meaningful improvement over Pix2Act’s score of 46.7, with a margin of 15.6. Although this result does not yet reach human-level performance, it demonstrates the potential of our approach to solve real-world computer tasks with minimal human effort in data collection.
@@ 5.3 Ablation studies
CoT inducing instructions for CAAP prompting
@@
Figure 4 presents experiment results comparing the effects of human demonstrations and CoT inducing instructions in our CAAP prompting method. It is shown that the use of CoT inducing instructions results in a meaningful improvement in performance. Removing CoT inducing instructions from the prompt results in a 2.3% drop compared to the baseline, and a 3.8% drop when the human demo is absent. When only rationales are excluded from human demonstrations, there is a reduction in task-solving SR by 3.5% relative to the fully-prompted baseline. Excluding the whole human demonstrations leads to a decrease of 9.5%. Additionally, the removal of both CoT-inducing instructions and human demonstrations results in an decrease in SR by as much as 13.3%, demonstrating that effective ICL provided by our CAAP prompting scheme significantly enhances the decision-making capacity of the LLM model in use. The per-task result are depicted in Table 10 of Appendix J.1.
@@
Since our main experiments combine GUI-based perception and execution with the CAAP prompting technique, we conducted additional experiments to isolate the effect of CAAP prompting as part of an ablation study. Without modifying the input/output channels, we tested how the performance of the RCI agent (Kim et al., 2023) improves when its prompting mechanism is replaced by our CAAP prompting. We first executed the original RCI code from the authors‘ official GitHub repository without modifications, testing only the 47 tasks listed in the repository by 50 times and taking the average. ***The success rate of the RCI agent we observed deviates from the results reported in the literature, despite our diligent efforts to replicate the experimental setup exactly. Upon inquiry, the RCI authors indicated that they no longer had access to the complete experimental logs used in their publication. Then, we replaced the prompting method with CAAP and reran the tests. To ensure a fair comparison, we used identical seeds for task instances, the same demo set provided by the repository, and the same LLM for queries.
As described in Table 3, the results demonstrate that applying CAAP prompting to the RCI agent significantly increased the average task-solving SR by 7.1%, from 76.9% to 84.0%. (Task-level results can be seen in Table 11 of Appendix J.2.) The prompting methods of RCI and CAAP exhibit two key differences. First, the RCI approach tends to modify even correct action plans during the iterative critique and revision process. In contrast, the CAAP method, by focusing on thorough context review rather than critique, is able to avoid this issue when formulating action plans. Second, RCI executes actions sequentially based on a fixed initial plan, failing to adapt to changes in the environment that occur during execution. Conversely, the CAAP method is designed to account for state changes resulting from prior actions. This experiment underscores the effectiveness of our CAAP prompting, confirming its potential to enhance the performance of LLMs in computer tasks.
@@ 6 Conclusion
In this paper, we introduce a modular agent designed to mimic how humans solve tasks on computers. Specifically, we aim to address the limitations of existing models that rely on imitation learning with large quantities of human demonstrations or non-universal input sources such as HTML and DOM. Our agent operates solely on visual inputs and significantly reduces the need for human demonstrations by leveraging the common knowledge and reasoning capabilities built into LLMs. Additionally, we introduce the CAAP prompting mechanism to enhance the decision-making capabilities of LLM-based agents on computer tasks. Evaluations on the MiniWoB++ and WebShop benchmarks demonstrate that our agent outperforms previous methods while using significantly less data. By adopting a modular approach that separates visual perception, action planning, and action execution, our method allows for easy adaptation to new tasks or model upgrades. Additionally, our agent is designed to operate beyond web environments and across inter-application settings, offering broader applicability and significant advantages in terms of generality.
@@
Appendix A MiniWoB++ Experiment Details
Models for visual observer
Fine-tuning dataset for visual observer
Data collection for fine-tuning the visual observer involves a two-step process: screenshot collection and annotation of the collected UI elements. Initially, the target task is manually performed while screenshots are taken at each step. Image comparisons are then conducted to remove duplicate screenshots. From 10 episodes for each of the 73 tasks, a total of 1,768 screenshots were collected. In the second stage, a GUI-based annotation tool, as depicted in Figure 6, is utilized. Each UI element within the screenshots is marked with bounding boxes, and specific attributes are assigned. These attributes are subsequently stored in a JSON format. Data augmentation techniques are also employed to enhance the dataset. These techniques include color adjustments, minor scaling of bounding box sizes, and the random insertion of Gaussian noise. As a result, the total number of annotated data items is tripled compared to the original dataset.
The YOLOv8 model is fine-tuned using the screenshots and their bounding boxes around the UI elements. The Pix2Struct model is fine-tuned using the masked screenshots (Fig. 2(a)) and the corresponding element attributes, excluding the positions of bounding boxes.
Human demo dataset
Train and test episode split
Appendix B Comparisons to other approaches
Method Modality Model Training Experiment
Vision Decision Scheme MiniWoB++ WebShop
Human - - - - -
[…]
WebGUM HTML+Image ViT-B16 Flan-T5-XL SL $\surd square-root\surd$
[…]
- / Image only - -
Pix2Act Image Pix2Struct Transformer SL+RL $\surd square-root\surd$
[…]
CAAP Image
YOLOv8,
Pix2Struct
gpt-4-0125 SL $\surd square-root\surd$
B.1 Experiment result for MiniWoB++
The data requirements and reported performance in literature are described in Table 5. One notable approach is CC-Net, which reports results across the broadest range of MiniWoB++ tasks. However, this required the collection of 2.4 million demonstrations, demanding approximately 6,300 hours of human labor, and faces challenges in real-world applications due to its high reliance on DOM information. RCI (Kim et al., 2023) and AdaPlanner (Sun et al., 2023) achieved average task SRs above 92% with fewer than 100 demonstrations, but their dependence on HTML inputs limits their task coverage. When recalculating SRs, including unsupported or unreported tasks, their performance drops to 50.8% and 49.2%, respectively. WebGUM (Furuta et al., 2023) expanded task coverage by using both HTML and image inputs. However, despite utilizing 401K demonstrations, its results only showed a 1.9% improvement over RCI, achieving an SR of 52.7%. Pix2Act (Shaw et al., 2023) was the first method to achieve high task SRs on MiniWoB++ using image inputs exclusively. It leveraged 1.3 million demonstrations, achieving over 96% success on 58 tasks, with an SR of 55.8% across the same set of 100 tasks. In contrast, SeeClick, which also used image inputs but on a smaller demonstration dataset, failed to sufficiently demonstrate problem-solving ability. Our approach represents the most advanced image-only agent, solving 73 tasks with a high SR of 94.5%, using fewer than 1.8K demonstrations. Furthermore, it achieved a 63.3% SR on the 100-task range, 7.5% higher than Pix2Act, with 61 tasks demonstrating an SR of over 80%—five more tasks than Pix2Act. (See Table 12 of Appendix J.3.) This represents a significant improvement achieved with fewer than 100 human demonstrations, resulting from the powerful reasoning capabilities of the LLM and our carefully modularized agent design.
Human - - - 93.5% 104
[…]
WebN-T5 HTML 0 12K 48.4% 56
[…]
WebGUM HTML+Image 0 401K 94.2% 56
$C$-Net DOM+Image 0 2.4M 93.6% 104
$C$-Net Image 0 2.4M 24.0% 104
Pix2Act Image 0 1.3M 96.2% 59
SeeClick Image 0 2.8K 69.4% 55
CAAP Image 1.76K 0.1K 94.5% 73
B.2 Experiment result for WebShop
Table 6 compares our approach with studies reporting performance on the WebShop benchmark. WebGUM was trained using both HTML and image inputs with 1K human demonstrations, achieving an average reward of 66.6, surpassing the performance of the initial model introduced with the WebShop benchmark, which was trained using an SL+RL approach. Pix2Act, designed to operate solely on image inputs, reported an average reward of 46.7 when trained with 1K human demonstrations using a combination of SL and RL techniques. Our agent, utilizing a vision model training dataset collected through an automated process and just a single human demonstration, achieved a significantly higher average reward of 62.3, outperforming Pix2Act. This result is comparable to methods that leverage both HTML and image inputs, demonstrating that our approach can effectively function across web and non-web environments with minimal human effort in dataset collection.
Table 6: Comparison with other methods that solve WebShop problems.
Method Modality Image datasize Demo datasize Task score SR
Human - - - 82.1 59.6%
WebShop HTML+Image 0 10.6K 62.4 28.7%
WebGUM HTML+Image 0 1K 67.5 45.0%
Pix2Act Image 0 1K 46.7 N/A
CAAP Image 0.3K 1 62.3 34.7%
Appendix C Element coverage of CAAP agent
@@
Appendix D Action coverage of CAAP agent
@@
Appendix E Additional related work
Vision models for agents
Numerous advanced techniques have been developed for detecting and interpreting UI elements on screens. Faster R-CNN (Ren et al., 2015) employs a Region Proposal Network (RPN) to extract areas of interest, followed by accurate object detection within these regions. This approach offers high accuracy in complex scenes, though at the cost of slower processing speed. The Single Shot MultiBox Detector (SSD) (Liu et al., 2016) offers rapid inference by eliminating the RPN and maintains decent accuracy. To improve small object detection and reduce false positives, one can diversify default boxes and feature maps for the SSD. Co-DETR (Zong et al., 2023), which leverages a Transformer-based architecture, performs detection while learning the relationships between objects. This anchor-free approach simplifies the learning process and effectively understands object interrelations, but it requires significant time and resources for training. YOLO (Redmon et al., 2016) processes predictions for object classes and bounding boxes across the entire image in one go. YOLO has both high speed and high accuracy compared to most other models. Unlike the previously mentioned technologies that primarily focus on extracting positional and typological information of objects, Pix2Struct (Lee et al., 2023) specializes in understanding detected objects. Designed around a Vision Transformer (ViT)(Dosovitskiy et al., 2020) and a T5 transformer(Raffel et al., 2020), it takes an image as input and outputs structured text, playing a crucial role in extracting and analyzing detailed attributes and embedded text information. Our approach leverages the combination of YOLO and Pix2Struct for visual information processing, aiming to capitalize on the strengths of both high-speed object detection and detailed attribute analysis.
Appendix F Case study of MiniWoB++ tasks
In this section, we closely examine the failure cases that arise during the execution of MiniWoB++ tasks by the CAAP agent. Figure 5 presents screenshots of selected exemplary failure cases. We have categorized these eight instances into three groups: incorrect directives, observation failures, and decision-making failures.
Figure 5: Fail cases of the CAAP agent.
Incorrect directives
Observation failures
Decision making failures
In Figure 5, cases (e), (f), (g), and (h) illustrate failures resulting from decision-making errors. Case (e) exemplifies a task requiring the identification of the sixth word, ’sed.’ Although the visual observer correctly observed and reported the scene, the LLM incorrectly identified the sixth word. In case (f), the task was to find ’Cassandry’, which was not visible on the screen. The LLM initiated a progressive scroll but prematurely stopped scrolling and failed to make the correct decision to continue. Case (g) involved the LLM determining the coordinates of text-written blocks and predicting their positions to move an object labeled ’Hermione’ to the central left position. The LLM decided on the drag action but incorrectly positioned the endpoint, leading to failure. Case (h) required highlighting a specific paragraph within a document composed of multiple paragraphs. Our visual observer is designed to detect text on a line-by-line basis, making it a challenging task to predict paragraph structure from the coordinates of each line. In this case, our agent failed to accurately predict the paragraph configuration. These four cases represent failures due to incorrect judgments by the LLM, with particular difficulty, as seen in (g) and (h), in interpreting coordinates represented numerically.
Among the failure cases, a tiny portion of the cases result from errors in the benchmark directives. While roughly 40% arise due to observation errors inherent in the agent configuration based on visual observation data, the others are failures that occur within the decision making domain. However, most minor errors in directives or observations tend to be appropriately accounted for during the decision making process by the LLM, often leading to successful outcomes as much as failures. This tendency is expected to intensify as the reasoning capabilities of LLMs advance, thereby improving their task-solving abilities.
Appendix G Annotation tool
Appendix H Prompt examples
@@
Figure 8: An example of a CAAP prompt.
@@
Figure 9: An example of LLM response to the CAAP prompt.
Appendix I Function calling example
@@
Appendix J Task-level result of MiniWoB++ Experiments
J.1 Per-task results for ablation study on CAAP components
Task Full Demo without CoT phrases CoT phrases with No CoT,
configurations CoT phrases without demo action-only demo no demo
choose-list 1.000 0.980 0.940 0.920 0.980
circle-center 1.000 1.000 1.000 0.980 0.980
click-button 0.980 0.980 0.980 0.980 0.980
click-button-sequence 0.920 0.920 0.920 0.920 0.920
click-checkboxes 0.980 1.000 1.000 1.000 0.960
click-checkboxes-large 1.000 0.960 1.000 1.000 0.860
click-checkboxes-soft 1.000 0.980 0.980 1.000 0.960
click-checkboxes-transfer 1.000 1.000 1.000 1.000 1.000
click-collapsible 0.920 0.140 1.000 0.960 0.340
click-collapsible-2 0.900 0.900 0.920 0.920 0.780
click-dialog 1.000 1.000 1.000 1.000 1.000
click-dialog-2 1.000 1.000 0.980 1.000 1.000
click-link 1.000 1.000 1.000 1.000 1.000
click-menu 0.800 0.880 0.360 0.400 0.360
click-menu-2 0.920 1.000 0.560 0.660 0.780
click-option 0.960 0.960 0.960 0.980 0.940
click-scroll-list 0.940 0.940 0.860 0.920 0.700
click-tab 1.000 1.000 1.000 1.000 1.000
click-tab-2 0.960 0.980 0.980 0.960 0.980
click-tab-2-easy 1.000 1.000 1.000 1.000 1.000
click-tab-2-hard 0.980 1.000 1.000 1.000 1.000
click-tab-2-medium 1.000 1.000 1.000 1.000 1.000
click-test 1.000 1.000 1.000 1.000 1.000
click-test-2 0.920 0.920 0.920 0.920 0.920
click-test-transfer 0.920 0.920 0.920 0.920 0.920
click-widget 1.000 0.960 0.860 0.860 0.860
copy-paste 0.980 0.980 0.980 0.980 0.960
copy-paste-2 1.000 1.000 0.820 0.980 0.780
drag-box 1.000 1.000 0.980 1.000 0.960
drag-circle 0.980 1.000 0.960 0.980 1.000
drag-items 0.880 0.860 0.940 0.940 0.580
drag-items-grid 0.780 0.620 0.600 0.780 0.620
drag-single-shape 0.940 0.980 0.980 0.940 0.980
drag-sort-numbers 0.720 0.480 0.560 0.620 0.160
email-inbox 0.960 0.960 0.860 0.900 0.720
email-inbox-delete 1.000 1.000 1.000 1.000 0.980
email-inbox-forward 0.960 0.980 0.640 0.960 0.620
email-inbox-forward-nl 0.960 1.000 0.460 0.920 0.520
email-inbox-forward-nl-turk 0.820 0.800 0.220 0.820 0.380
email-inbox-important 1.000 1.000 0.980 1.000 0.960
email-inbox-nl-turk 0.900 0.800 0.540 0.860 0.500
Task Full Demo without CoT phrases CoT phrases with No CoT,
configurations CoT phrases without demo action-only demo no demo
email-inbox-noscroll 0.980 0.940 0.920 1.000 0.860
email-inbox-reply 0.960 0.980 0.940 0.900 0.960
email-inbox-star-reply 0.980 0.960 0.980 0.960 0.960
enter-password 1.000 1.000 1.000 1.000 0.980
enter-text 1.000 1.000 1.000 1.000 0.980
enter-text-2 1.000 1.000 1.000 1.000 0.980
enter-text-dynamic 1.000 1.000 1.000 1.000 0.980
find-greatest 1.000 1.000 0.420 0.860 1.000
find-word 0.820 0.620 0.740 0.800 0.380
focus-text 1.000 1.000 1.000 1.000 0.980
focus-text-2 1.000 1.000 1.000 1.000 0.980
generate-number 0.860 0.780 0.840 0.880 0.680
guess-number 0.980 1.000 0.860 0.960 0.900
highlight-text 0.920 0.960 0.680 0.900 0.940
highlight-text-2 0.460 0.500 0.140 0.400 0.240
identify-shape 0.980 0.980 0.940 0.960 0.960
login-user 1.000 1.000 1.000 1.000 0.980
login-user-popup 1.000 0.820 0.440 0.440 0.440
multi-layouts 0.980 0.920 0.960 1.000 0.660
multi-orderings 1.000 1.000 0.960 1.000 0.660
odd-or-even 0.980 0.920 0.980 0.920 0.940
read-table 0.940 0.820 0.900 0.900 0.840
read-table-2 0.740 0.700 0.720 0.720 0.700
resize-textarea 1.000 1.000 0.900 0.980 0.980
scroll-text 0.620 0.780 0.200 0.420 0.020
scroll-text-2 1.000 0.980 0.720 0.880 0.780
simple-algebra 1.000 0.980 1.000 1.000 0.700
simple-arithmetic 1.000 1.000 1.000 1.000 0.940
sign-agreement 0.960 0.980 0.420 0.780 0.420
unicode-test 0.900 0.920 0.900 0.880 0.880
use-autocomplete 0.960 0.880 0.900 0.960 0.900
use-spinner 0.980 1.000 0.900 0.920 0.700
Average SR 0.945 0.922 0.850 0.910 0.812
J.2 Comparison of the CAAP and RCI promptings
Table 11: Per-task SRs for RCI agent with CAAP and original RCI agent.
Task RCI RCI
agent agent
with (original)
CAAP
choose-list 0.80 1.00
click-button 1.00 1.00
click-button-sequence 1.00 1.00
click-checkboxes 1.00 1.00
click-checkboxes-large 1.00 1.00
click-checkboxes-soft 0.98 0.84
click-checkboxes-transfer 1.00 1.00
click-collapsible 1.00 1.00
click-collapsible-2 0.92 0.82
click-color 1.00 1.00
click-dialog 1.00 1.00
click-dialog-2 1.00 1.00
click-link 0.98 1.00
click-menu 1.00 1.00
click-option 1.00 1.00
click-scroll-list 1.00 1.00
click-shades 1.00 1.00
click-shape 0.94 0.94
click-tab 1.00 1.00
click-tab-2 0.92 0.94
click-tab-2-hard 0.98 0.96
click-test 1.00 1.00
click-test-2 0.98 1.00
click-widget 1.00 1.00
count-shape 0.78 0.76
Task RCI RCI
agent agent
with (original)
CAAP
email-inbox 0.66 0.82
email-inbox-forward-nl 0.90 0.02
email-inbox-forward-nl-turk 1.00 0.20
email-inbox-nl-turk 0.40 0.28
enter-date 0.00 0.00
enter-password 1.00 1.00
enter-time 0.02 0.00
focus-text 0.98 1.00
grid-coordinate 0.96 1.00
identify-shape 1.00 0.76
login-user-popup 0.44 0.44
multi-layouts 0.88 0.82
navigate-tree 0.92 0.86
search-engine 0.30 0.30
simple-algebra 0.80 0.40
social-media 0.98 0.96
social-media-all 0.98 0.98
social-media-some 0.90 0.44
terminal 0.00 0.00
tic-tac-toe 0.30 0.14
use-autocomplete 0.92 0.48
use-spinner 0.88 1.00
Average SR 0.840 0.769
J.3 Comparison of the CAAP agent with SoTAs in per-task result

### v3

@@ 5.2 WebShop
WebShop (Yao et al., 2022b) is a simulated e-commerce web environment well-suited for assessing an agent’s ability to handle real-world scenarios. To successfully complete WebShop tasks, the agent must understand textual instructions and requirements for the target product and navigate various types of web pages to locate it, out of a pool of 1.18 million real products scraped from amazon.com. The environment provides a total of 12,087 instructions, divided into 10,587 for training, 1,000 for development, and 500 for testing. Among them, we utilized only the 500 test instructions for evaluation purposes.
Models and datasets
The web browser for the WebShop environment can be resized freely, and for convenience, we used a fixed resolution of 1276 × 1153 pixels for both data collection and evaluation. Unlike MiniWoB++, the dataset used to train the visual observer for WebShop was generated entirely through an automated process. Given an instruction, a simple helper agent we created extracted a search term, performed a search, and clicked the first item in the results. Screen images were captured throughout this process, and annotation data was collected using Selenium for each image, resulting in a total of 300 annotated samples. The action-proposal model, supported action types, and function-calling approach are entirely consistent with those employed in the MiniWoB++ experiments. For the evaluation of the CAAP agent, a single human demonstration is applied universally to all instructions.
Experiment result
There has been limited research on constructing agents that rely solely on image inputs, and Pix2Act is the only method whose performance has been validated on the WebShop benchmark. Table 2 presents a comparison between our approach and Pix2Act on WebShop. (Comparisons with other agent-based methods that rely on additional HTML input, on top of image input, are presented in Appendix I.2.) While Pix2Act employed 1,012 human demonstrations to train its agent model, our approach relied on just a single human demonstration, along with 300 samples collected through a fully automated process for training the visual observer module. Despite the significantly reduced manual data collection effort, our agent achieved a task score of 62.3, a meaningful improvement over Pix2Act’s score of 46.7, with a margin of 15.6. Although this result does not yet reach human-level performance, it demonstrates the potential of our approach to solve real-world computer tasks with minimal human effort in data collection.
@@ 5.3 Ablation studies
Impact of CoT inducement on task-solving performance
@@
Figure 4 presents experimental results illustrating the impact of CoT inducement on task performance. The results show that incorporating CoT-inducing instructions leads to a significant improvement in performance. Specifically, as shown in the figure, adding CoT-inducing instructions increases task SRs by 5.3% (from 79.2% to 84.5%) in the absence of human demonstrations and by 2.5% (from 92.0% to 94.5%) when human demonstrations are included. Moreover, augmenting human demonstrations with rationales further enhances task SRs by 3.5%, raising them from 91.0% to 94.5%. Additionally, removing human demonstrations entirely results in a 10.0% drop in task SRs, while removing CoT-inducing instructions in addition leads to a substantial decrease of 15.3%. These findings highlight that CAAP prompting, by explicitly and implicitly guiding the LLM to think step-by-step and establish contextual relationships before proposing actions, significantly improves task SRs in solving computer-based tasks. Detailed per-task results are presented in Table 11 of Appendix J.
@@
Since our main experiments combine GUI-based perception and execution with the CAAP prompting technique, we conducted additional experiments to isolate the effect of CAAP prompting as part of an ablation study. Without modifying the input/output channels, we tested how the performance of the RCI agent (Kim et al., 2023) improves when its prompting mechanism is replaced by our CAAP prompting. We first executed the original RCI code from the authors’ official GitHub repository without modifications, testing only the 47 tasks listed in the repository by 50 times and taking the average.***The success rate of the RCI agent we observed deviates from the results reported in the literature, despite our diligent efforts to replicate the experimental setup exactly. Upon inquiry, the RCI authors indicated that they no longer had access to the complete experimental logs used in their publication. Then, we replaced the prompting method with CAAP and reran the tests. To ensure a fair comparison, we used identical seeds for task instances, the same demo set provided by the repository, and the same LLM for queries.
As described in Table 3, the results demonstrate that applying CAAP prompting to the RCI agent significantly increased the average task-solving SR by 7.1%, from 76.9% to 84.0%. (Task-level results can be seen in Table 12 of Appendix K.) The prompting methods of RCI and CAAP exhibit two key differences. First, the RCI approach tends to modify even correct action plans during the iterative critique and revision process. In contrast, the CAAP method, by focusing on thorough context review rather than critique, is able to avoid this issue when formulating action plans. Second, RCI executes actions sequentially based on a fixed initial plan, failing to adapt to changes in the environment that occur during execution. Conversely, the CAAP method is designed to account for state changes resulting from prior actions. This experiment underscores the effectiveness of our CAAP prompting, confirming its potential to enhance the performance of LLMs in computer tasks.
@@ 6 Conclusion
In this paper, we introduce a modularly designed GUI agent model that emulates the way humans solve tasks on computers. Our modular design not only effectively enhances pixel coordinate detection—a critical challenge faced by integrated VLM architectures—but also addresses the limitations of existing models that rely heavily on imitation learning with large amounts of human demonstrations. Furthermore, we propose the CAAP prompting mechanism to improve the action-proposal capabilities of LLM-based agents for computer tasks by guiding the LLM to fully utilize the CoT reasoning mechanism. Evaluations on the MiniWoB++ and WebShop benchmarks demonstrate that our agent model achieves superior performance compared to the previous methods while requiring significantly less data. By adopting a modular approach that separates visual perception, action proposal, and action execution, our method allows for easy adaptation to new tasks or model upgrades. Additionally, our agent is designed to operate beyond web environments and across inter-application settings, offering broader applicability and significant advantages in terms of generality.
@@
Appendix A Element coverage of CAAP agent
@@
Appendix B Action coverage of CAAP agent
@@
Appendix C MiniWoB++ Experiment Details
Models for visual observer
Fine-tuning dataset for visual observer
Data collection for fine-tuning the visual observer involves a two-step process: screenshot collection and annotation of the collected UI elements. Initially, the target task is manually performed while screenshots are taken at each step. Image comparisons are then conducted to remove duplicate screenshots. From 10 episodes for each of the 73 tasks, a total of 1,768 screenshots were collected. In the second stage, a GUI-based annotation tool, as depicted in Figure 10 of Appendix H, is utilized. Each UI element within the screenshots is marked with bounding boxes, and specific attributes are assigned. These attributes are subsequently stored in a JSON format. Data augmentation techniques are also employed to enhance the dataset. These techniques include color adjustments, minor scaling of bounding box sizes, and the random insertion of Gaussian noise. As a result, the total number of annotated data items is tripled compared to the original dataset.
The YOLOv8 model is fine-tuned using the screenshots and their bounding boxes around the UI elements. The Pix2Struct model is fine-tuned using the masked screenshots (Figure 2(a)) and the corresponding element attributes, excluding the positions of bounding boxes.
Human demo dataset
Train and test episode split
Appendix D Prompt examples
@@
Figure 6: An example of the CAAP prompt. The raionales contained in the expert demonstrations and the CoT-inducing instructions, which are detailed in the ’### Instructions ###’ section, help an LLM gradually develop the semantic connectivity between the task description, the action history, and the visual state of the screen, in a CoT manner.
@@
Figure 7: An example of an LLM response to the CAAP prompt in Figure 6, showing how the LLM improves evidence for action-proposal by connecting the surrounding context information.
Appendix E Function calling example
@@
Appendix F Task coverage of the agents that rely solely on screen images for MiniWoB++
Appendix G Comparison of the CAAP agent with SoTAs in MiniWoB++ per-task result